# DBudgetKV: Dynamic Budget in KV Cache Compression for Ensuring Optimal Performance

### v1

@@ 3.3 Implemetation Details
Additionally, we have discovered that applying DBudgetKV to the model’s 0th and 1st layers results in poor performance when the model operates under a high budget. We hypothesize that this may be due to the relatively uniform distribution of attention scores in these initial layers, leading to the premature discarding of tokens that are crucial for subsequent layers, ultimately causing the entire model’s inference to collapse. A more detailed analysis of this issue is provided in Appendix B. Based on this finding, we initiate the KV cache compression operations starting from the model’s 2nd layer.
@@ 4.1 Experimental Settings
For comprehensive evaluation, we evaluate DBudgetKV using datasets of both short and long context length. For tasks of relatively short inputs, the models are assessed across mathametics, sicence, and commonsense reasoning on multiple datasets including GSM8K Cobbe et al. (2021), GPQA Rein et al. (2023), TheoremQA Chen et al. (2023), TruthfulQA Lin et al. (2022), and CoQA Reddy et al. (2019). For long-context tasks, we adopt tasks and datasets from Longbench Bai et al. (2024). Appendix C provides a detailed description and statistics of these datasets along with how they are utilized in experiments.
@@
Our primary objective is to evaluate the capability of DBudgetKV to compress KV cache while striking for a full-cache performance. To this end, we compare DBudgetKV with the full KV cache performance and calculate the average compression ratio per dataset. Additionally, we select three previous KV cache compression methods with fixed budget size, including: Heavy Hitter Oracle (H2O) Zhang et al. (2023), StreamingLLM (SLM) Xiao et al. (2024), and SnapKV Li et al. (2024). We compare with two budgets for these three methods at 90% and 50% compression ratio respectively.
@@ 4.2 Main Results
• DBudgetKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types. Interestingly, with Llama3-8B and Qwen2.5-7B, DBudgetKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache. With Mistral, DBudgetKV also manages to achieve nearly 20% KV cache compression with only a 1.5% performance reduction. These results indicate that our proposed method can be practically deployed in real-world scenarios without worrying domain-specific budget thresholds. In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g. 90%. However, when the budget is reduced, e.g. 50%, the degradation can become severe on certain datasets, distinct from DBudgetKV that automatically adjusts the pruning to always attain full performance.
Methods GSM8K CoQA NQA Musique QMSum
@@
Table 2: Performance of LLMs with scales of 70B, 34B, and 72B across five datasets using DBudgetKV.
@@
Table 3: Performance comparison with different $k$ values, ranking methods and universal thresholds.
• DBudgetKV natually reflects the difficulty of the generation task. As in Table 1, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%). This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation.
Methods GSM8K CoQA NarrativeQA Musique QMSum TriviaQA Avg.
Llama3-8B-Instruct
Full 4.693 0.289 3.441 3.659 5.458 2.717 3.376
Llama3-70B-Instruct
Full 17.504 1.167 6.285 6.565 13.993 5.302 8.469
4.3 DBudgetKV’s Generalization Ability
We next conduct experiments to examine whether our design and hyperparameters can be generalized to LLMs of larger sizes. On five datasets in Table 2, DBudgetKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same DBudgetKV setting. Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% lossless compression at its peak, validating the robustness and versatility for its general applicability.
4.4 Ablation Study
The ablation study is conducted to explore the impact of various configurations of DBudgetKV. We select Llama-3-8B-Instruct and perform experiments across five datasets presented in Table 3.
Attention Matrix Reduction
The reduced attention matrix $A^{\prime}$ in Sec. 3.2 aggregates attention scores from the last $k$ positions. The upper part of Table 3 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ (n being the number of input tokens). It is evident that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$. On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget). What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last token, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length. The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
Importance Ranking Strategies
DBudgetKV simply regards positions as the significance of tokens, while another intuitive strategy is to utilize attention scores to rank their importance. We investigate its potentials through ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O. The results are shown in the middle of Table 3: under the same experimental settings, the attention-based ranking struggles to maintain the model’s performance. Though it achieves a high compression ratio, the performance is not stable with possible severe degradation, which deviates from our objective. Therefore, the position-based strategy is more suited for our goal.
Attention Norm Threshold
As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in the bottom part of Table 3. Intuitively, a larger threshold allows for a higher compression ratio, while a smaller threshold does the opposite. The conclusion is clear that when the threshold is set to a smaller 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement. Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates. Thus, we deem 1% as a reasonable threshold for universal uses across models and tasks.
Appendix D presents additional results and analysis from the ablation study.
4.5 Efficiency Analysis
@@ 5 Conclusion
In this study, we introduce an innovative KV cache compression objective designed to approximate the full-cache performance, independent of specific inputs, while optimizing resource utilization through targeted KV cache pruning. Our approach, termed DBudgetKV, employs a straightforward yet effective two-step process for each layer of LLMs. Tokens are initially ranked solely based on their positions, followed by the eviction of the least significant tokens as determined by the norm value of the reduced attention matrix. Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that DBudgetKV achieves nearly lossless compression, while with notable space and time reduction, successfully fulfilling our objective for greater practical values.
Limitations
Appendix A Full Algorithm of DBudgetKV
Algorithm 1 DBudgetKV
@@ Appendix A Full Algorithm of DBudgetKV
Appendix B Freeze the first two layers of LLM
In this section, we elaborate the necessity to freeze the first and second layers in DBudgetKV as mentioned in Section 3. In our preliminary studies, we apply DBudgetKV to Llama2-7B-Chat and conduct experiments on GSM8K and CoQA. We first perform a case study, followed by a comparison with the effects of not freezing the bottom layers.
As shown in Table 5, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence. However, the budget exceeds 90% in both cases. By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 0 and 1 are relatively low, while the budgets from layer 2 to layer 30 become very high, with the last layer being low again. We hypothesize that the model is still encoding the global semantics in its early layers, thus not yet able to identify truly important tokens in the first two layers, of which the attention distribution is relatively uniform. This phenomenon has also been observed by early works on Transformers analysis Ethayarajh (2019); Garí Soler and Apidianaki (2021). Not freezing early layers may lead to early eviction of important tokens, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
Based on the above case study, we attempt to freeze certain layers of the model and explore the optimal freezing configuration. We continue to validate the results of freezing different layers, and the resuls are shown in Table 6. We can observe that freezing the first two layers achieves a balance between model performance and budget. Moreover, freezing the last layer does not significantly enhance the model’s performance and instead leads to an increase in budget. DBudgetKV ultimately opts to begin KV cache compression from the 2nd layer of the model.
@@ Appendix B Freeze the first two layers of LLM
Table 5: Case Study.
Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
Table 6: Performance of Llama2-7B-Chat on GSM8K and CoQA using DBudgetKV with different frozen layers.
Appendix C Datasets Used in Experiments
Math & Science
Commonsense Reasoning (CR)
Single Document QA (Single-Doc QA)
This task assesses the model’s reading comprehension skills when dealing with a single, extended document. NarrativeQA Kociský et al. (2018) is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes. Qasper Dasigi et al. (2021) is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information. We employ F1 Score as the metric for above two datasets.
Multi-Document QA (Multi-Doc QA)
This task evaluates the model’s reading comprehension capabilities across multiple extended documents. 2WikiMultiHopQA (2WKMQA) Ho et al. (2020) is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles. MuSiQue Trivedi et al. (2022) evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately. We leverage F1 Score as the metric for above two datasets.
Methods GSM8K CoQA NQA Musique QMSum
Full 75.28 52.74 24.06 14.77 22.27
Post=1% Performance Using Different $k$
$k=1$ 76.50 52.86 23.44 15.96 21.95
Budget 93.2% 81.5% 48.7% 43.7% 15.0%
$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
Budget 95.1% 94.8% 77.3% 77.2% 59.5%
$k=2\%n$ 76.19 52.82 22.87 14.34 21.74
Budget 96.0% 96.1% 69.8% 62.0% 47.1%
$k=3\%n$ 75.82 52.80 23.03 13.61 21.56
Budget 98.6% 95.7% 60.3% 46.0% 39.5%
$k=4\%n$ 75.21 52.76 22.96 13.45 21.47
Budget 98.8% 95.8% 51.9% 35.2% 35.4%
$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
Budget 98.6% 96.2% 45.1% 30.0% 33.1%
$k=6\%n$ 75.66 52.81 21.95 14.33 20.94
Budget 98.5% 96.6% 40.4% 25.8% 31.6%
$k=7\%n$ 76.57 52.88 21.18 13.66 20.94
Budget 98.0% 96.7% 37.3% 22.5% 30.6%
$k=8\%n$ 76.35 52.86 20.36 13.75 21.11
Budget 97.4% 96.7% 35.1% 20.9% 30.1%
$k=9\%n$ 76.65 52.84 19.99 13.66 20.76
Budget 96.8% 96.8% 33.4% 19.9% 29.5%
$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
Budget 96.3% 96.8% 32.0% 19.7% 29.3%
Table 7: Performance comparison with different $k$values, ranking methods and universal thresholds.
Summarization
This task examines the model’s ability to comprehend and summarize lengthy documents. QMSum Zhong et al. (2021) is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions. Multi-News (M-News) Fabbri et al. (2019) is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic. We use Rouge-L as the metric for above two datasets.
Few-Shot Learning (FSL)
This task assesses the model’s few-shot learning capabilities. TriviaQA Joshi et al. (2017) is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall. We use F1 Score as the metric.
Code
This task evaluates the model’s ability to complete and generate code. LCC Guo et al. (2023) is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures. We use Edit Sim as the metric.
Appendix D Ablation
In this section, we present additional ablation study results. By setting various values for $k$, we expand upon the results shown in Table 3. The expanded results are displayed in Table 7. These experiments facilitate a deeper understanding of how different parameter settings impact model performance and provide a basis for optimizing parameter selection.

### v2

@@ 3.3 Implemetation Details
Additionally, we have discovered that applying DBudgetKV to the model’s 0th and 1st layers results in poor performance when the model operates under a high budget. We hypothesize that this may be due to the relatively uniform distribution of attention scores in these initial layers, leading to the premature discarding of tokens that are crucial for subsequent layers, ultimately causing the entire model’s inference to collapse. A more detailed analysis of this issue is provided in Appendix C. Based on this finding, we initiate the KV cache compression operations starting from the model’s 2nd layer. Regarding ablation studies are further conducted in Sec.4.5.
@@ 4.1 Experimental Settings
For comprehensive evaluation, we evaluate DBudgetKV using datasets of both short and long context length. For tasks of relatively short inputs, the models are assessed across mathematics, science, and commonsense reasoning on multiple datasets including GSM8K [14], GPQA [26], TheoremQA [27], TruthfulQA [28], and CoQA [29]. For long-context tasks, we adopt tasks and datasets from Longbench [30]. Appendix A provides a detailed description and statistics of these datasets along with how they are utilized in experiments.
@@
Our primary objective is to evaluate the capability of DBudgetKV to compress KV cache while striking for a full-cache performance. To this end, we compare DBudgetKV with the full KV cache performance and calculate the average compression ratio per dataset. Additionally, we select three previous KV cache compression methods with fixed budget size, including: Heavy Hitter Oracle (H2O) [8], StreamingLLM (SLM) [10], and SnapKV [18]. We compare with three budgets for these three methods at 90%, 50% and 20% compression ratio respectively.
@@ 4.2 Main Results
Table 3: Performance of LLMs with scales of 70B, 32B, and 72B across five datasets using DBudgetKV.
GSM8K CoQA NQA Musique QMSum
@@
• DBudgetKV is able to fulfill our objective, capable of performing near-lossless dynamic compression across different models, varying input lengths, and diverse task types. Interestingly, with Llama3-8B and Qwen2.5-7B, DBudgetKV even surpasses the full-cache performance by 0.12% and 2.63% respectively, utilizing an average of 63.68% and 76.02% KV cache. With Mistral, DBudgetKV also manages to achieve nearly 20% KV cache compression with only a 1.5% performance reduction. These results indicate that our proposed method can be suited for real-world scenarios with no bother by domain-specific budget thresholds. In stark contrast, previous methods achieve a consistent full-cache performance only when manually determined a high budget ratio, e.g. 90%. However, when the budget is reduced, e.g. 50% and 20%, the degradation can become severe on certain datasets, distinct from DBudgetKV that automatically adjusts the pruning to always prioritize full performance.
• DBudgetKV naturally reflects the difficulty of the generation task. As in Table 2, the dynamic budget ratio is high on Math&Science datasets (over 90%), while much lower on QA or Summarization datasets (as low as 15%). This observation is in line with our intuition, where inference on concise but hard tasks, such as math problems, requires more context and more precise calculation, resulting in higher budget allocation. Thus, our method design serves beyond for inference efficiency, but could also be leveraged for input analysis in a broader scope.
It is worth noting that the scope of this work is not to propose yet another KV cache pruning method. Instead, we identify the common limitation of previous KV pruning methods, and call for optimal KV pruning without a pre-defined budget. Consequently, the consistent performance across domains after KV pruning is the ultimate goal of our method, instead of achieving the highest compression ratio on certain datasets. Under our objective, we leave the further improvement on compression ratio as a promising future research direction.
4.3 DBudgetKV’s Generalization Ability
We next conduct experiments to examine whether our design and hyper-parameters can be generalized to LLMs of larger sizes. On five datasets in Table 3, DBudgetKV with Llama3-70B and Qwen2.5-32B/72B demonstrates consistent near full-cache performance with the same DBudgetKV setting. Notably, the averaged compression ratio increases on datasets of longer context, achieving nearly a 50% compression. But most importantly, they are able to maintain lossless performance across inputs, validating the robustness and versatility for its general applicability.
4.4 Efficiency Analysis
GSM8K CoQA NarrativeQA Musique QMSum TriviaQA Avg.
Llama3-8B-Instruct
Full 4.693 0.289 3.441 3.659 5.458 2.717 3.376
Llama3-70B-Instruct
Full 17.504 1.167 6.285 6.565 13.993 5.302 8.469
Table 5: Performance comparison for ablation studies in Section 4.5, with different $k$ values for attention matrix reduction, ranking methods and universal thresholds.
@@ 4.4 Efficiency Analysis
4.5 Ablation Study
The ablation study is conducted to explore the impact of various configurations of DBudgetKV. We select Llama-3-8B-Instruct and perform experiments across five datasets presented in Table 5. Appendix B presents additional results and analysis from the ablation study.
Attention Matrix Reduction
The reduced attention matrix $A^{\prime}$ in Sec. 3.2 aggregates attention scores from the last $k$ positions. The upper part of Table 5 illustrates the model’s performance and the actual budget when setting $k=1$, $1\%n$, $5\%n$, and $10\%n$ ($n$ being the number of input tokens). It is clear that setting $k$ as $1$ achieves a significantly reduced budget, thus a higher compression ratio, with almost no change in performance compared to $1\%n$ and $5\%n$. On the other hand, while $10\%n$ can compress more KV cache, it fails to maintain performance (for instance, on NarrativeQA, the former achieves a performance of 9.74 using 32% of the budget, whereas the latter scores 21.44 using 48.7% of the budget). What’s even better is that since $k=1$ requires the least amount of computation, relying solely on the scores from the last token, the complexity of obtaining $A^{\prime}$ becomes $O(1)$, independent of the sequence length. The advantages of both high efficacy and low overhead make $k=1$ a solid design choice for calculating the norm metric.
Importance Ranking Strategies
We investigate DBudgetKV potentials through ranking by each token’s average attention score received from other tokens, similar to previous approaches such as H2O. The results are shown in the middle of Table 5: under the same experimental settings, the attention-based ranking struggles to maintain the model’s performance. Therefore, the position-based strategy is more suited for our goal.
Table 6: Performance comparison between DBudgetKV and StreamingLLM with the first two layers frozen (SLMF) under difference KV cache budgets ($0.9$, $0.5$ and $0.2$).
GSM8K GPQA CoQA NrtvQA QMSum TriviaQA
Full 75.28 29.02 52.74 24.06 22.27 70.26
Oursk=1 76.50 30.13 52.86 23.44 21.95 64.24
Budget 93.2% 86.7% 81.5% 48.7% 15.0% 41.0%
SLM0.9 72.78 28.79 52.83 23.75 22.48 69.97
SLMF0.9 76.42 29.24 52.89 22.95 22.47 69.10
SLM0.5 3.41 6.03 44.68 20.94 21.80 67.79
SLMF0.5 4.55 3.13 44.59 21.31 21.88 68.04
SLM0.2 3.21 3.35 40.11 20.12 21.11 66.17
SLMF0.2 1.21 1.12 38.44 19.52 20.61 65.69
Attention Norm Threshold
As we adopt the universal threshold as 1% for the attention norm difference, we comprehensively study the effects of smaller or larger thresholds, as shown in the bottom part of Table 5. Intuitively, a larger threshold allows for a higher compression ratio, while a smaller threshold does the opposite. The conclusion is clear that when the threshold is set to a smaller value, such as 0.1%, the average budget increases as expected, yet the model’s performance sees little improvement. Conversely, when the threshold is set to 10%, it prunes more KV cache, but the model’s performance significantly deteriorates. Thus, we deem 1% as a reasonable threshold for universal uses across models and tasks.
Comparison with StreamingLLM
We perform additional experiments on six datasets using StreamingLLM (with Llama3-8B) that has the first two layers frozen, as an ablation study on the effects of the proposed pruning and halting process. As shown in Table 6, despite having the first two layers frozen, StreamingLLM still falls short of achieving the objective proposed in this work, not able to remain optimal consistently across datasets like GSM8K and GPQA as the budget varies. It is demonstrated that solely leveraging the position bias as in StreamingLLM is not able to maintain the full performance; the proposed pruning and halting mechanism that integrates the attention norm metric is necessary in achieving the goal for lossless pruning.
@@ 5 Conclusion
In this study, we introduce a novel KV cache compression objective designed to approximate the full-cache performance, independent of specific inputs, while optimizing resource utilization through targeted KV cache pruning. Our approach, termed DBudgetKV, employs a straightforward yet effective two-step process for each layer of LLMs. Tokens are initially ranked solely based on their positions, followed by the eviction of the least significant tokens as determined by the norm value of the reduced attention matrix. Comprehensive experiments conducted across diverse datasets, encompassing a variety of tasks and context lengths, demonstrate that DBudgetKV achieves nearly lossless compression, while with notable space and time reduction, successfully fulfilling our objective for greater practical values.
Appendix A Datasets Used in Experiments
Math & Science
Commonsense Reasoning (CR)
Single Document QA (Single-Doc QA)
This task assesses the model’s reading comprehension skills when dealing with a single, extended document. NarrativeQA [31] is a dataset designed to evaluate model’s ability to comprehend and answer questions based on narrative texts, focusing on understanding stories and their underlying themes. Qasper [32] is a dataset aimed at assessing model’s capability to extract and answer questions from academic papers, emphasizing understanding complex scientific information. We employ F1 Score as the metric for above two datasets.
Multi-Document QA (Multi-Doc QA)
This task evaluates the model’s reading comprehension capabilities across multiple extended documents. 2WikiMultiHopQA (2WKMQA) [33] is a dataset designed to test model’s ability to perform multi-hop reasoning and answer complex questions using information from multiple Wikipedia articles. MuSiQue [34] evaluates model’s skill in integrating and reasoning over information from multiple sources to answer comprehensive questions accurately. We leverage F1 Score as the metric for above two datasets.
Summarization
This task examines the model’s ability to comprehend and summarize lengthy documents. QMSum [35] is a dataset for evaluating model’s ability to generate concise summaries of meeting transcripts, focusing on capturing the key points from multi-party discussions. Multi-News (M-News) [36] is a dataset that challenges models to create coherent summaries by synthesizing information from multiple news articles on the same topic. We use Rouge-L as the metric for above two datasets.
Few-Shot Learning (FSL)
This task assesses the model’s few-shot learning capabilities. TriviaQA [37] is a dataset designed to assess model’s ability to retrieve and answer questions based on large collections of trivia, emphasizing comprehension and factual recall. We use F1 Score as the metric.
Code
This task evaluates the model’s ability to complete and generate code. LCC [38] is a dataset focused on evaluating models’ ability to understand and generate code by considering extended code contexts, enhancing the ability to reason over complex programming structures. We use Edit Sim as the metric.
Appendix B Ablation
Table 7: Performance comparison with different $k$, expanded from Table 5.
Methods GSM8K CoQA NQA Musique QMSum
Full 75.28 52.74 24.06 14.77 22.27
Post=1% Performance Using Different $k$
$k=1$ 76.50 52.86 23.44 15.96 21.95
Budget 93.2% 81.5% 48.7% 43.7% 15.0%
$k=1\%n$ 76.19 52.87 23.17 15.40 21.94
Budget 95.1% 94.8% 77.3% 77.2% 59.5%
$k=2\%n$ 76.19 52.82 22.87 14.34 21.74
Budget 96.0% 96.1% 69.8% 62.0% 47.1%
$k=3\%n$ 75.82 52.80 23.03 13.61 21.56
Budget 98.6% 95.7% 60.3% 46.0% 39.5%
$k=4\%n$ 75.21 52.76 22.96 13.45 21.47
Budget 98.8% 95.8% 51.9% 35.2% 35.4%
$k=5\%n$ 75.59 52.75 21.62 13.71 21.43
Budget 98.6% 96.2% 45.1% 30.0% 33.1%
$k=6\%n$ 75.66 52.81 21.95 14.33 20.94
Budget 98.5% 96.6% 40.4% 25.8% 31.6%
$k=7\%n$ 76.57 52.88 21.18 13.66 20.94
Budget 98.0% 96.7% 37.3% 22.5% 30.6%
$k=8\%n$ 76.35 52.86 20.36 13.75 21.11
Budget 97.4% 96.7% 35.1% 20.9% 30.1%
$k=9\%n$ 76.65 52.84 19.99 13.66 20.76
Budget 96.8% 96.8% 33.4% 19.9% 29.5%
$k=10\%n$ 76.72 52.85 9.74 13.67 21.11
Budget 96.3% 96.8% 32.0% 19.7% 29.3%
In this section, we present additional ablation study results for Section 4.5. By setting various values for $k$, we expand upon the results shown in Table 5. The expanded results are displayed in Table 7. These experiments facilitate a deeper understanding of how different parameter settings impact model performance and provide a basis for optimizing parameter selection.
Appendix C Freeze the first two layers of LLM
In this section, we demonstrate the necessity of using DBudgetKV to compress the KV cache starting from the 2nd layer of LLMs. We apply DBudgetKV on Llama2-7B-Chat and conduct experiments on GSM8K and CoQA. We first perform a case study, followed by a comparison of the effects of not freezing layers of the model on its performance.
As shown in Table 9, when the threshold is set to 1% and no layers are frozen, the outputs for two examples on GSM8K and CoQA are incorrect and lack logical coherence. However, the budget exceeds 90% in both cases. By observing the actual budget of each layer, we can see that for these two examples, the budgets for layer 0 and 1 are relatively low, while the budgets from layer 2 to layer 30 are the same and very high, with the last layer being low again. We analyze that since the model, in the first two layers, is not yet able to identify truly important tokens, the attention distribution is relatively uniform. This may lead to the model discarding truly important tokens during eviction, resulting in subsequent generations being unable to access this information, ultimately causing the output to fail.
Table 8: Performance on GSM8K and CoQA using DBudgetKV with different frozen layers.
Datasets Layer None 0 0,1 0,1,2 0,1,2,3 0,1,31
GSM8K Oursk=1 0.014 0.250 0.264 0.252 0.264 0.252
Budget 93.1% 97.0% 95.4% 97.5% 97.7% 98.2%
CoQA Oursk=1 1.47 52.80 53.58 53.58 53.32 53.46
Budget 92.7% 94.7% 95.5% 98.3% 98.4% 99.2%
Based on the above case study, we attempt to freeze certain layers of the model and explore the optimal freezing configuration. We continue to test the results of freezing different layers. The resuls are shown in Table 8. From the table, we can observe that freezing the first two layers achieves a balance between model performance and budget. Moreover, freezing the 31st layer does not significantly enhance the model’s performance and instead leads to an increase in budget. Consequently, DBudgetKV ultimately opts to begin KV cache compression from the 2nd layer of the model.
Appendix D Full Algorithm of DBudgetKV
@@
Algorithm 1 DBudgetKV
Table 9: Case Study.
@@
Appendix E Limitations
Appendix F Broader Impacts
Our work has made contributions on both a new objective in this line of research, as well as a new method that successfully fulfills this objective. Since all current KV cache pruning methods require a pre-defined KV budget threshold, they either have inconsistent and unstable performance across different input lengths, domains and tasks, or they need careful tuning of their budgets per dataset to ensure the full-cache performance. Our proposed method evaluated on 13 diverse datasets, including short and long-context QA, coding and math, demonstrates its effectiveness on addressing this limitation, which automatically adjusts the budget to reach full-cache performance regardless of the inputs.